Instrument Log

MPL 02           C1                  CF Lamont, OK Optical Trailer

Date

Operator

Comments

8/14/98

Mentor

Instrument removed by mentor. Replaced with high resolution MPL. See MPLHR logs.

8/14/98

ZW

The MPL is down at this time. The mentor is on site, and working on this instrument.

7/16/98

MW

The MPL monitor was displaying a garbled screen this morning. Could not read the data or view the graphs. Rebooted the computer at 1540 GMT. RTS at 1544 GMT.

7/1/98

MW

The MPL display was found to be garbled this morning. Could not read any data from the display. Rebooted the computer at 1558 GMT. The display returned to normal. RTS at 1600 GMT.

6/16/98

SS

Work order 2900 stated that the SDS reports no files due collection failed. Upon arriving on site discovered the display was scrambled. Quit the program and shut down the computer. Rebooted the computer at 1855. System appears normal at 1857. RTS at 1900 GMT.

5/26/98

MK

Due to a power failure on 5/25/98 at 0425 GMT, the MPL monitor was not displaying any data, and the emission light was not on. Powered on the emission at 1451 GMT. Status OK. RTS at 1506 GMT.

5/20/1998

SS

During daily checks observed the Laser beam lens to be dirty. Secured the Laser beam output at 1450 GMT. Cleaned the Laser beam lens. Restored the Laser beam output at 1452 GMT.

5/15/98

SS

Due to power failure restarted instrument at 1525 GMT.

5/7/98

SS

No beam output observed. Pressed emission to start beam emission. RTS at 1328 GMT.

5/5/98

MW

The MPL was not emitting a beam and no data was being collected. The 'Emission' LED was not lit. Pressed the 'Emission' button and the beam came on. Exited the MPL.BAT program, shut down the computer and rebooted at 1512 GMT. RTS at 1515 GMT.

4/2/98

MW

The MPL02 monitor was garbled this morning. Could not read any data or the graph. Rebooted the computer at 1517 GMT. RTS at 1520 GMT. Status OK.

3/23/98

SS

Observed water spots on the laser beam transmitter output. Secured the laser emission at 1435 GMT. Cleaned the laser beam lens. Returned the laser to emission at 1436 GMT. RTS at 1445 GMT.

3/18/98

SS

Found system with no AC power applied. Found circuit breaker on UPS tripped. Reset at 1515 GMT. System RTS at 1543 GMT.

3/17/98

SS

While checking instrument the diode power tripped from pulse to external. Reset diode power to pulse. Returned to service at 1515 GMT.

3/16/98

MW

The MPL02 beam was not transmitting. The pulse LED and the Emission light were not lit. Depressed the Pulse button and Emission button at 2015 GMT. The MPL02 returned to emission mode. RTS. Suspect power bump caused the laser to stop transmitting.

3/5/98

SS

During daily checks found display screen garbled and unreadable. Exited program and performed a system reboot. Screen is now readable and has good data. RTS at 1545 GMT.

2/18/98

KT

The screen was cluttered so system was rebooted at 1631 GMT. RTS at 1646 GMT.

1/5/98

MW

The digital readout on the spectra-physics keypad indicated 0.02 Watts of output power. The emission lights on the laser diode module and the keypad were not lit. No beam was observed and the graph displayed no cloud data. Pressed the "control" button on the keypad at 1546 GMT. Beam turned on and cloud data resumed immediately after pressing the button.

12/10/97

KT

At 1555 GMT the MPL was power cycled due to the monitor screen being scattered. Status O.K. RTS at 1559 GMT.

11/25/97

CAM

SDS TR indicated no files had been received since 11/24/97 at 2202 GMT. Check of instrument found that the display was all jumbled with white vertical lines through the screen and the data was not clearly visible. Rebooted system at 1442 GMT. RTS status O.K. at 1447 GMT.

11/10/97

GRS

The normal display window was not up. The desk top was up with an OS2 window open. Went ahead and performed a Monday morning reboot of the system at 1600 GMT. System came on line at 1614 GMT. RTS at 1615 GMT.

11/4/97

MW

The MPL display was garbled this morning. Typed "ctrl-esc" and selected "shut down" at 1548 GMT. Pressed "ctrl-alt-del" at 1549 GMT. RTS at 1551 GMT.

10/20/97

GRS

The data seemed reasonable but there were some (abnormal) lines outside of the data window. Performed a reboot of the system from 1600 to 1615 GMT. RTS at 1617 GMT.

10/6/97

MW

The MPL02 has been dismantled by the mentor so some calibrations can be performed. The MPL02 should be back online this afternoon. No corrective maintenance performed at this time.

9/15/97

MW

The MPL02 is down for repair by the instrument mentor. No corrective maintenance performed.

9/12/97

MW

The instrument mentor has taken the MPL off-line to do some maintenance. No T/S performed by site ops.

9/8/97

MW

The MPL was not emitting a beam and the physics keypad showed 0.3 watts. Also, the emission light, pulse light, and watts light were not lit on the keypad. The system light was on. The monitor display did not look normal. There were horizontal bars across the screen. Performed a reboot of the system computer at 1538 GMT. Pressed the pulse light button on the keypad and the pulse led lit up. Pressed the emission button on the keypad and the emission led lit up. Pressed the system button and the watts light came on. The displayed wattage on the physics keypad went to 1.0 watts. However, there was still no beam being emitted. Decreased the wattage on the keypad to 0.96 watts and then increased it back up to 1.0 watts. The beam turned on. RTS at 1545 GMT.

9/3/97

MDW

Due to a Kay Electric transformer going down in Billings on 9/2, the site lost all power at around 2030 GMT. The extended power outage caused the optical trailer UPS to drain its batteries and all instruments lost power. When power was restored on 9/3 at 0142 GMT the MPL02 laser did not come up in the emission mode and the MPL.bat program did not restart. Returned the Spectra-Physics keypad to emission mode at 1402 GMT on 9/3. Rebooted computer at 1412 GMT. System RTS status O.K. at 1415 GMT.

9/2/97

MW

Discovered the MPL02 screen was garbled this morning during the daily PM rounds. The graphs and data were completely unreadable. Performed a reboot at 1530 GMT. "MPL.bat" did not start as it should. Only the "FTP.exe" file started. Performed a second reboot at 1540 GMT. "Telnet.exe" and "FTP.exe" started but "MPL.bat" did not start. Right-clicked on the screen and selected "Find." Selected the button labeled "Search all subgroups" and typed "MPL.bat" in the file box. The "MPL.bat" icon appeared. Started "MPL.bat" by double-clicking on the icon at 1448 GMT. The normal MPL01 program came up and began collecting data.

7/28/97

CM/CW

AC power to the optical trailer had to be shut off at 1845 GMT for construction. Performed graceful shut down at 1830 GMT. AC power restored at 1905 GMT and system RTS status O.K. at 1905 GMT.

7/8/97

MW

The MPL data acquisition window was all garbled. It was impossible to read anything on the screen. Rebooted the MPL computer at 1420 GMT. RTS at 1428 GMT.

6/23/97

ZW

Performed weekly reboot of system.

6/16/97

MK

Performed weekly reboot of system.

6/2/97

MR

Performed weekly reboot of system.

5/19/97

ZW

Performed weekly reboot of MPL at 1455 GMT.

5/12/97

ZW

Performed weekly reboot of MPL at 1440 GMT.

4/28/97

MS

Worked with Conner Flynn over the telephone to get MPL back online. The problem of not re-booting correctly was corrected by moving a file associated with CRON_2 program. RTS 1600 GMT.

4/25/97

GRS

While performing the evening shutdown at the Optical trailer it was noticed that the data did not look reasonable and that the filter detector temperature read 0. Performed a graceful shutdown at 2054 GMT and power cycled the computer. The computer stopped in the middle of its boot up with a CRON2.EXE error. Performed two more graceful shutdowns and power cycles with the same error on boot up. Called the Mentor at 2110 GMT and left a voice message. Conner Flynn called back at 2130 GMT. The problem was described to him and since it was past closing time we decided to wait until Monday morning to try and fix. The problem is quite probably a software error. T/S in progress.

4/14/97

Mentor

Connor Flynn on site and replaced the entire system with a new one. System is now MPL 04.

4/11/97

MW

The MPL window assembly had condensation built up on the inside of the window and around the tube wall. This condensation seemed to be scattering the beam so you could not determine the cloud height on the display. Placed the cover over the beam at 1456 GMT. Removed the window assembly and cleaned all of the condensation out. Replaced the assembly and removed the beam cover at 1511 GMT. RTS.

4/11/97

MW

The MPL has been taking data in 10 second intervals since 4/4/97 at 2100 GMT when Connor Flynn requested the change for troubleshooting purposes. Rebooted the system at 1722 GMT on 4/11/97. System now taking data at 1 min intervals. RTS at 1730 GMT.

3/10/97

Mentor

Connor Flynn on site this week for maintenance/upgrades. Data will be missing and/or corrupted intermittently during this time.

2/17/97

MR

Rebooted MPL at 1458 GMT.

2/10/97

MR

Cleaned the laser window at 1455 GMT. Rebooted MPL at 1520 GMT. RTS at 1523 GMT.

1/28/97

KDT

Daily check. Performed weekly reboot of the system. Status OK.

1/21/97

DKB

Daily check. Performed system re-boot at 1453 GMT. Had some software errors but was able to get started at 1520 GMT.

1/09/97

CAM

Significant ice/snow build up on window. Thawed as much as possible by hand. Remaining ice left to thaw naturally.

12/30/96

CAM

Weekly reboot of the system at 1517 GMT.

12/17/96

CAM

Laser was not emitting. T/S indicated the Spectra Physics keypad was not in pulsed emission mode. Returned keypad to pulsed emission mode and RTS at 1945 GMT.

12/16/96

CAM

Weekly reboot of the system at 1516 GMT. Computer detected a system error (cron2.exe) after reboot. Ended program and RTS at 1512 GMT.

12/02/96

CAM

Weekly reboot of the system at 1509 GMT. Computer detected a system error (cron2.exe) after reboot. Ended program and RTS at 1512 GMT.

11/25/96

CAM

Weekly reboot of the system at 1500 GMT.

11/12/96

CAM

Daily PM. Weekly re-boot of the system at 1500 GMT.

11/4/96

CAM

Daily PM. Weekly re-boot of the system at 1515 GMT.

10/28/96

CAM

Daily PM. Weekly re-boot of the system at 1339 GMT.

10/21/96

CAM

Daily PM. Weekly re-boot of the system at 1635 GMT.

10/4/96

GRS

Daily PM. Found program locked up. Re-booted and tested OK at 1504 GMT.

9/12/96

CAM

Daily PM. Found program locked up since 9/11 1906 GMT. Re-booted and RTS at 1616 GMT.

9/11/96

CAM

Daily PM. Cleaned window at 1436 GMT.

9/4/96

CAM

Daily PM. Cleaned window at 1415 GMT.

8/29/96

GRS

Daily PM. Found display garbled. Re-booted and RTS at 1433 GMT. Data lost from 8/28 2000 to 8/29 1433 GMT.

8/19/96

CAM

Daily PM. Cleaned window at 1457 GMT.

8/09/96

GRS

Program locked up at 2359 GMT on 8/08/96. Rebooted computer at 1439 GMT. RTS status O.K. at 1440 GMT.

7/24/96

CAM

1339 GMT Daily PM. Cleaned windows.

7/16/96

GRS

1417 GMT System locked up with same error as above. Called mentor and left voice message of problem. Rebooted computer at 1417 GMT and system RTS at 1420 GMT.

6/12/96

CAM

Daily check found program stopped with "internal processing" error at location ## 0160:fff5fbd5 -000d:9bd5.60000,9084. Re-booted and RTS at 1336 GMT.

6/7/96

CAM

Daily check cleaned window at 1415 GMT.

5/16/96

GRS

Daily check cleaned dirt from window at 1400 GMT. Found data display garbled. Stopped program, re-booted, and RTS at 1413 GMT.

4/15/96

Mentor

R/R *optical laser. * Old - ? * New - 1528702

4/11/96

CAM

Daily check found display blank and program stopped. Re-booted and RTS at 1444 GMT.

3/26/96

CAM/GRS

Daily check. Monitor displaying error message <2:11:41>c:\>. Informed mentor and re-booted at 1456 GMT. SDS did not collect any files the first hour. Performed another re-boot and RTS at 1606 GMT.

3/8/96

Mentor

R/R *Laser Diode, **power supply and ***optical laser. Installed parts from system 01 and tested OK. * Old - 3515 * New - 1096491 ** Old - 0316 ** New - ? *** Old - 1528702 *** New - ?

2/28/96

Mentor

R/R *Laser Diode, **power supply and ***optical laser. * Old - 1528703 * New - 3515 ** Old - ? ** New - 0316 *** Old - ? *** New - 1528702

2/1/96

Mentor

System is inop awaiting parts.

1/30/96

Mentor

2000 GMT. Powered down to move to new location.

1/29/96

TMG/CAM

AC power secured for construction at 1423 GMT. RTS at 2045 GMT.

1/16/96

TMG/CAM

1449 GMT. Daily checks. Found the program locked up since 1/12 2200 GMT. Re-booted and RTS at 1452 GMT.

12/27/95

TMG/CAM

AC power secured for construction at 1458 GMT.